# RadOnc-GPT: An Autonomous LLM Agent for Real-Time Patient Outcomes Labeling at Scale

### v1

@@ 2.7 Adjudication Procedure
For the tier 2 (complex) outcomes tasks, discrepancies between RadOnc-GPT’s results and baseline labels were independently adjudicated by oncologists with domain expertise specific to each clinical scenario (prostate cancer, head-and-neck cancer, mandibular osteoradionecrosis), one oncologist per dataset. Oncologists classified discrepancies explicitly into three categories:
@@ 3.1 Tier 1 Task - Structured Data Extraction (QA)
In the initial structured extraction QA tier, RadOnc-GPT demonstrated exceptionally high fidelity in accurately retrieving structured demographic and treatment-related information. Across all 500 evaluated patients (totaling 3,000 individual demographic fields), RadOnc-GPT matched the demographic database records perfectly (100%). Additionally, RadOnc-GPT accurately reproduced the treatment planning information in 497 out of 500 cases (99.4%). These results established foundational trust in RadOnc-GPT’s basic retrieval capability, crucial for subsequent complex clinical outcomes labeling based on structured and unstructured patient data.
@@ 3.2 Tier 2 Tasks - Complex Clinical Outcomes Labeling based on Structured and Unstructured patient data
Mandibular Osteoradionecrosis (ORN) in Head-and-Neck Patients:
Among 233 head-and-neck patients evaluated for ORN, RadOnc-GPT initially achieved an accuracy of 84.5%, with high recall (88.2%) but lower precision (48.4%). After adjudication by an independent radiation oncologist, overall accuracy significantly improved to 95.2%. The adjudication process revealed that of the 36 initial discrepancies (32 false positives, 4 false negatives):
@@
Head-and-Neck Cancer Recurrence:
For the head-and-neck cancer recurrence cohort (82 patients), initial accuracy was 92.7%, with a precision of 92.0% and recall of 95.8%. After independent adjudication, accuracy significantly increased to 96.3%. The 6 initial discrepancies (4 false positives, 2 false negatives) were adjudicated as follows:
@@
Across three complex patient outcomes labeling tasks (ORN in head–and–neck, prostate cancer recurrence, head–and–neck cancer recurrence; total $n=395$ before adjudication), expert review of all discordant cases ($n=48$) found 30 (63%) baseline labeling errors, 13 model errors, and 5 indeterminate cases. Post–adjudication, false negatives fell from 9 to 4, and pooled (micro–averaged) performance improved from Precision/Recall/F1/Accuracy of 74.3/92.6/82.5/87.8% to 90.3/97.0/93.6/95.4% (Table 3). Note that post–adjudication denominators exclude indeterminate cases for the affected patient cohorts.
Table 3: Confusion counts and performance metrics before and after adjudication. Totals are pooled (micro–averaged) across tasks. Post–adjudication NN excludes indeterminate cases (ORN: 4; H&N recurrence: 1).
@@ 4 Discussion
This study investigated whether an autonomous, self-retrieving LLM agent can simultaneously curate patient outcomes and audit patient data quality within a realistic clinical setting. We adopted a two-tier evaluation strategy by design: first establishing fidelity in structured data retrieval before advancing to the more complex task of patient outcome labeling. In Tier 1, RadOnc-GPT successfully reproduced structured data alone, achieving exact matches for all demographic fields and accurately reproducing per-course treatment information in 497 of 500 patients, including course identifiers, ICD codes, delivered plan identifiers, and radiation modality. Because these comparisons were made against structured database values from Aria, no manual review was required, making this approach well-suited for routine quality assurance. Importantly, this QA test also surfaced a systematic error early in our evaluation: course IDs with leading numbers were occasionally truncated by the LLM. Identifying this issue allowed us to implement corrections and strengthen the retrieval pathway. Thus, Tier 1 not only served as a validity checkpoint but also provided a mechanism for hardening the retrieval process before progressing to more complex evaluative tasks.
@@
For Tier 2 tasks, after adjudication, recall reached 100% for ORN detection, 92.5% for prostate cancer recurrence detection, and 97.9% for head-and-neck cancer recurrence detection. In the context of surveillance and registry curation, false negatives are typically unobserved whereas false positives remain reviewable; thus, a recall-oriented operating point is preferable. The ORN finding—absence of false negatives—is particularly noteworthy, as missed late toxicities carry substantial clinical consequences. Two design choices likely contributed to this performance: (1) prompts explicitly requiring evidence both for and against each Marx stage, and (2) function-call–guided retrieval with task-specific temporal bounds, which maintains clinical context while filtering out irrelevant text. Together, these elements mitigate risks such as fragmentation of clinically linked facts across retrieval chunks and the limitations of generic embeddings, which often underperform in radiotherapy-specific domains.
@@
Adjudication reframed model-versus-label disagreements. Of 48 initial discrepancies, 30 (63%) were attributable to baseline labeling errors, 13 to model-generated errors, and 5 remained indeterminate. After excluding indeterminate cases, the adjusted accuracies of the original institutional labels were 95.2% for ORN, 97.5% for prostate cancer recurrence, and 96.3% for head-and-neck cancer recurrence. Sutton et al. reported for head-and-neck cancer a registry accuracy of 89.4% with sensitivity 61% and specificity 99%, further noting that longer time-to-recurrence was associated with false negatives (median 1,007 vs. 414 days for true positives) [34]. Against this backdrop, RadOnc-GPT achieved higher recall while maintaining high accuracy, and it surfaced labeling errors that would otherwise have persisted. Although the comparison is indirect—owing to differences in definitions and cohorts—the findings suggest that autonomous agents can enhance completeness relative to conventional registry processes, while simultaneously generating an audit trail to support adjudication.
@@
Throughput supports near-real-time use. With four-way parallelism and per-patient latencies of 10-20 seconds, most cohorts can be processed within a single day. Outputs are returned as structured JSON, enabling seamless integration into dashboards, analytics pipelines, and trial-matching platforms. This operational profile is important because registry accuracy is time-sensitive: delays in documentation contribute to missed events in conventional workflows [34]. A nightly or on-demand run can substantially shorten this gap. Notably, the same cancer recurrence detection prompt was applied to both prostate and head-and-neck cancer, suggesting that disease-site generalization is feasible with foundational LLM-based systems.
@@
Limitations of this study include its reliance on single-center data and single-reviewer adjudication for each dataset; multi-reviewer panels could shift borderline determinations. The underlying model is proprietary (GPT-4o), and both governance considerations and financial cost may constrain adoption, with continuous institution-wide runs remaining cost-sensitive at present. These factors motivate future benchmarking against open-source models, exploration of incremental recomputation and caching strategies, and deployment patterns that respect privacy requirements (including federated approaches). Finally, although the agent generated rationale-rich outputs that facilitated adjudication, we did not evaluate human–AI conflict resolution policies in live operations; such mechanisms will be essential for routine deployment.
@@ 5 Conclusions
Because the approach is prompt-driven, requires no model fine-tuning, and leverages granular yet generic API hooks, it is readily extensible to other oncology centers and disease sites. We envision autonomous LLM agents functioning as continuous auditors—running nightly against new patients, updating outcomes registries in real time, and allowing clinicians to focus on judgment rather than data wrangling. The anticipated result is richer and timelier evidence to inform precision radiotherapy, ultimately advancing patient care.
Acknowledgements
Institutional policy required that any physician with greater than 21 patients involved be included as an author.
Disclosures

### v2

@@ 2.7 Adjudication Procedure
For these complex outcomes labeling tasks (tier 2), discrepancies between RadOnc-GPT’s results and baseline labels were independently adjudicated by oncologists with domain expertise specific to each clinical scenario (prostate cancer, HNC, mandibular osteoradionecrosis), one oncologist per dataset. Oncologists classified discrepancies explicitly into three categories:
@@ 3.1 Tier 1 Task - Structured Data Extraction (QA)
In the initial structured extraction QA tier, RadOnc-GPT demonstrated exceptionally high fidelity in accurately retrieving structured demographic and treatment-related information. Across all 500 evaluated patients (totaling 3,000 individual demographic fields), RadOnc-GPT matched the demographic database records perfectly (100%). Additionally, RadOnc-GPT accurately reproduced the treatment planning information in 497 out of 500 cases (99.4%). These results established foundational trust in RadOnc-GPT’s basic retrieval capability, a crucial prerequisite result before performing complex clinical outcomes labeling based on both structured and unstructured patient data.
@@ 3.2 Tier 2 Tasks - Complex Clinical Outcomes Labeling based on Structured and Unstructured patient data
Mandibular Osteoradionecrosis (ORN) in HNC Patients:
Among 233 HNC patients evaluated for ORN, RadOnc-GPT initially achieved an accuracy of 84.5%, with high recall (88.2%) but lower precision (48.4%). After adjudication by an independent radiation oncologist, overall accuracy significantly improved to 95.2%. The adjudication process revealed that of the 36 initial discrepancies (32 false positives, 4 false negatives):
@@
HNC Recurrence:
For the HNC recurrence cohort (82 patients), initial accuracy was 92.7%, with a precision of 92.0% and recall of 95.8%. After independent adjudication, accuracy significantly increased to 96.3%. The 6 initial discrepancies (4 false positives, 2 false negatives) were adjudicated as follows:
@@
Across three complex patient outcomes labeling tasks (ORN in HNC, prostate cancer recurrence, HNC recurrence; total $n=395$ before adjudication), expert review of all discordant cases ($n=48$) found 30 (63%) baseline labeling errors, 13 model errors, and 5 indeterminate cases. Post–adjudication, false negatives fell from 9 to 4, and pooled (micro–averaged) performance improved from Precision/Recall/F1/Accuracy of 74.3/92.6/82.5/87.8% to 90.3/97.0/93.6/95.4% (Table 3). Note that post–adjudication denominators exclude indeterminate cases for the affected patient cohorts.
Table 3: Confusion counts and performance metrics before and after adjudication. Totals are pooled (micro–averaged) across tasks. Post–adjudication NN excludes indeterminate cases (ORN: 4; HNC recurrence: 1).
@@ 4 Discussion
This study investigated whether an autonomous, self-retrieving LLM agent can simultaneously curate patient outcomes and audit patient data quality within a realistic clinical setting. We adopted a two-tier evaluation strategy by design: first establishing fidelity in structured data retrieval before advancing to the more complex task of patient outcome labeling. In Tier 1, RadOnc-GPT successfully reproduced structured data alone, achieving exact matches for all demographic fields and accurately reproducing per-course treatment information in 497 of 500 patients, including course identifiers, ICD codes, delivered plan identifiers, and radiation modality. Because these comparisons were made against structured database values from Aria (Varian Medical Systems, Palo Alto, CA), no manual review was required, making this approach well-suited for routine quality assurance. Importantly, this QA test also surfaced a systematic error early in our evaluation: course IDs with leading numbers were occasionally truncated by the LLM. Identifying this issue allowed us to implement corrections and strengthen the retrieval pathway. Thus, Tier 1 not only served as a validity checkpoint but also provided a mechanism for hardening the retrieval process before progressing to more complex evaluative tasks.
@@
For Tier 2 tasks, after adjudication, recall reached 100% for ORN detection, 92.5% for prostate cancer recurrence detection, and 97.9% for HNC recurrence detection. In the context of surveillance and registry curation, false negatives are typically unobserved since negatives are not typically reviewed, whereas false positives remain reviewable; thus, a recall-oriented operating point is preferable. The ORN finding—absence of false negatives—is particularly noteworthy, as missed late toxicities carry substantial clinical consequences. Two design choices likely contributed to this performance: (1) prompts explicitly requiring evidence both for and against each Marx stage, and (2) function-call–guided retrieval with task-specific temporal bounds, which maintains clinical context while filtering out irrelevant text. Together, these elements mitigate risks such as fragmentation of clinically linked facts across retrieval chunks and the limitations of generic embeddings, which often underperform in radiotherapy-specific domains.
@@
Adjudication reframed model-versus-label disagreements. Of 48 initial discrepancies, 30 (63%) were attributable to baseline labeling errors, 13 to model-generated errors, and 5 remained indeterminate. After excluding indeterminate cases, the adjusted accuracies of the original institutional labels were 95.2% for ORN, 97.5% for prostate cancer recurrence, and 96.3% for HNC recurrence. Sutton et al. reported a registry accuracy for HNC recurrence of 89.4% with sensitivity 61% and specificity 99%, further noting that longer time-to-recurrence was associated with false negatives (median 1,007 vs. 414 days for true positives) [SUTTON2024225]. Against this backdrop, RadOnc-GPT achieved higher recall while maintaining high accuracy, and it surfaced labeling errors that would otherwise have persisted. The findings suggest that autonomous agents can enhance completeness relative to conventional registry processes, while simultaneously generating an audit trail to support adjudication.
@@
Throughput supports near-real-time use. With rate-limited parallelism and per-patient latencies of 10-30 seconds, most cohorts can be processed within a single day. Outputs are returned as structured JSON, enabling seamless integration into dashboards, analytics pipelines, and trial-matching platforms. This operational profile is important because registry accuracy is time-sensitive: delays in documentation contribute to missed events in conventional workflows [SUTTON2024225]. A nightly or on-demand run can substantially shorten this gap. Notably, the same cancer recurrence detection prompt was applied to both prostate and HNC, suggesting that disease-site generalization is feasible with foundational LLM-based systems.
@@
Limitations of this study include its reliance on single-center data and single-reviewer adjudication for each dataset; multi-reviewer panels could shift borderline determinations. Additionally, although inference costs of commercial LLMs have been and will likely continue reducing over time, institution-wide runs using an agentic approach remain cost-sensitive at present. These factors motivate future benchmarking against open-source models, exploration of caching strategies, and prompt optimization of cheaper models. Finally, although the agent generated rationale-rich outputs that facilitated adjudication, we did not evaluate human–AI conflict resolution policies in live operations; such mechanisms will be essential for routine deployment.
@@ 5 Conclusions
Because the approach is prompt-driven and requires no model fine-tuning, it is readily extensible to other oncology centers and disease sites. We envision autonomous LLM agents functioning as continuous auditors, running nightly against new patients, updating outcomes registries in real time, and allowing clinicians to focus on patient care. The anticipated result is richer and timelier evidence to inform precision radiotherapy.
Author Contributions
J.H. created and developed RadOnc-GPT and conducted the experiments, Y.H. helped with writing and editing and formulating the study, V.C, L.M., and K.V.A. helped with reviewing and curating patient data and reviewing the manuscript, F.M., D.R., A.F., S.S., D.M., S.K., B.S., Y.G., M.N.W., S.V., and N.L. helped review the manuscript, S.P., M.F., and M.B. helped with testing and reviewing the manuscript, R.F. helped with reviewing ORN data, M.B.O. helped with curating prostate patient data, D.E., M.W., and S.R.B. helped review/compare RadOnc-GPT outputs, and W.L. was the P.I and co-creator.
Acknowledgement
This study was reviewed and approved by the Mayo Clinic Institutional Review Board (IRB; protocol number 24-010276). The requirement for informed consent was waived by the IRB due to the retrospective nature of the study. All procedures were conducted in accordance with the ethical standards of the Mayo Clinic and with relevant national and international guidelines and regulations.
Conflict of Interest
There are no conflicts of interest to report relating to this project.
Funding
This research was supported by NIH/BIBIB R01EB293388, by NIH/NCI R01CA280134, by the Eric & Wendy Schmidt Fund for AI Research & Innovation, and by the Kemper Marley Foundation.